OpenStudy (anonymous):

Show an equation and a solution for the problem. Arnold needs a 25% solution of nitric acid. He has 20 ml of a 30% solution. How many ml of a 15% solution should he add to obtain the required 25% solution? A. 20 ml B. 15 ml C. 25 ml D. 10 ml

OpenStudy (anonymous):

at the moment he has 30% of 20 which is \(.3\times 20=6\) ml of whatever is in the solution if he adds say \(x\) ml of 15% solution it will contribute \(.15x\) so the total of whatever is in the solution will be \(6+.15x\) on the other hand he will have \(20+x\) ml of solution all together which he wants to be \(25\%\) whatever so set \[6+.15x=.25(20+x)\] and solve for \(x\)

OpenStudy (anonymous):

probably easiest to get rid of the decimals by multiplying both sides by 100 and solve \[600+15x=25(20+x)\] etc